# Bridged Posterior: Optimization, Profile Likelihood and a New Approach of Generalized Bayes

### v1

@@ Abstract
Optimization is widely used in statistics, thanks to its efficiency for delivering point estimates on useful spaces, such as those satisfying low cardinality or combinatorial structure. To quantify uncertainty, Gibbs posterior exponentiates the negative loss function to form a posterior density. Nevertheless, Gibbs posteriors are supported in a high-dimensional space, and do not inherit the computational efficiency or constraint formulations from optimization. In this article, we explore a new generalized Bayes approach, viewing the likelihood as a function of data, parameters, and latent variables conditionally determined by an optimization sub-problem. Marginally, the latent variable given the data remains stochastic, and is characterized by its posterior distribution. This framework, coined “bridged posterior”, conforms to the Bayesian paradigm. Besides providing a novel generative model, we obtain a positively surprising theoretical finding that under mild conditions, the $\sqrt{n}$-adjusted posterior distribution of the parameters under our model converges to the same normal distribution as that of the canonical integrated posterior. Therefore, our result formally dispels a long-held belief that partial optimization of latent variables may lead to under-estimation of parameter uncertainty. We demonstrate the practical advantages of our approach under several settings, including maximum-margin classification, latent normal models, and harmonization of multiple networks.
KEY WORDS: Constrained conditional probability; Data augmentation; Duality; Profile likelihood; Parametric and semi-parametric Bernstein–von Mises.
@@ 1 Introduction
The Generalized Bayes approach is becoming increasingly popular due to its potential advantages in model simplicity and robustness. A generalized posterior can be specified based on partial information from the data, or via a loss function that characterizes an inferential summary of the data. This appeals when the likelihood is inaccessible or intractable; there is a well-established literature on partial information settings including methods based on composite likelihood (Lindsay, 1988; Varin et al., 2011), partial likelihood (Sinha et al., 2003; Dunson and Taylor, 2005), pairwise likelihood (Jensen and Künsch, 1994), and others. Recently, there has been a burgeoning interest in loss-based Bayesian models, including works involving classification loss (Polson and Scott, 2011) or distance-based losses (Duan and Dunson, 2021; Rigon et al., 2023; Natarajan et al., 2023). Loss-based generalized Bayes models typically use a probability distribution called the “Gibbs posterior” (Jiang and Tanner, 2008), taking the form:
$\Pi(\theta\mid y)\propto\exp\{-g(\theta,y)\},$
where $g(\theta,y)$ is some loss function with $y$ the data and $\theta$ the parameter.
@@
There is a vast generalized Bayes literature using the Gibbs posterior for explicit model weighting, with $g$ chosen according to some utility function such as predictive accuracy (Lavine et al., 2021; Tallman and West, 2023), scoring rule (Gneiting and Raftery, 2007; Dawid and Musio, 2015), fairness metrics (Chakraborty et al., 2024) or summary statistics-based divergence (Frazier and Drovandi, 2021; Frazier et al., 2023). Such an approach also lends itself to modular descriptions of data (Jacob et al., 2017), and can guard against model misspecification (Nott et al., 2023). With connections to these methods, our focus is on the cases when one wants to adopt a certain loss $g$ from the optimization literature for statistical modeling, while needing to quantify uncertainty beyond point estimates.
@@
In optimization, the point estimate $\hat{\theta}=\arg\min_{\theta}g(\theta,y)$ can often be efficiently computed using an iterative algorithm, even under a wide range of constraints. For example, convex clustering and its variants (Tan and Witten, 2015; Chi and Lange, 2015; Chakraborty and Xu, 2023) use $g(\theta,y)=(1/2)\sum_{i=1}^{n}\|y_{i}-\theta_{i}\|^{2}_{2}+\lambda\sum_{(i,j):i<j}\|\theta_{i}-\theta_{j}\|_{2}$ for data $y_{i}\in\mathbb{R}^{p}$, location parameter $\theta_{i}\in\mathbb{R}^{p}$, and tuning constant $\lambda>0$. This can be understood as a relaxation of hierarchical clustering; in place of a combinatorial constraint, the penalty term encourages most of the $L_{2}$-norms $\|\hat{\theta}_{i}-\hat{\theta}_{j}\|_{2}$ to be zero, promoting cluster structure via a small number of unique $\hat{\theta}_{i}$ at the solution. The estimate $\hat{\theta}$ can be obtained using convex, continuous optimization. A popular combinatorial alternative makes use of the $k$-means loss (MacQueen, 1967) toward clustering, $g\{(c_{1},\ldots,c_{n}),y\}=\sum_{k=1}^{K}\sum_{(i<j):c_{i}=c_{j}=k}\|y_{i}-y_{j}\|^{2}_{2}/n_{k}$, with $c_{i}\in\{1,\ldots,K\}$ where the discrete cluster assignment label $c_{i}=k$ if $\theta_{k}$ is the nearest centroid to $y_{i}$, $n_{k}=\sum_{i}1(c_{i}=k)$, and $\hat{\theta}_{i}=\sum_{i:c_{i}=k}y_{i}/n_{k}$. Here too, iterative algorithms can improve performance and better avoid local minima using continuous optimization techniques (Xu and Lange, 2019). Recently, Rigon et al. (2023) form a Gibbs posterior using this loss toward quantifying the uncertainty of $c_{i}$, which shows model robustness to the distributional asymmetry. Several recent works import long-standing ideas from optimization to account for constraints within a Bayesian framework (Duan et al., 2020; Presman and Xu, 2023; Zhou et al., 2024).
@@
Under the exponential negative transformation, the Gibbs posterior distribution concentrates near the posterior mode. This induces variability around the point estimate and, in turn, enables uncertainty quantification. How to interpret this uncertainty is not immediately obvious, and one may question the authenticity of inferential procedures such as hypotheses tests or intervals based on such a posterior which may not derive from a generative likelihood model. There are several recent works that provide sound justification for the large $n$ regime. First, Gibbs posteriors admit a coherent update scheme for $\theta$ toward minimizing the expected loss $\int g(\theta,y)\mathcal{F}(\textup{d}y;\theta)$, where $\mathcal{F}$ denotes the true data generating distribution (Bissiri et al., 2016). Second, if the Gibbs posterior density is proportional to a composite likelihood, such as the conditional density under some insufficient statistics (Lewis et al., 2021) derived from a full likelihood model $F(y;\theta_{0})$, then the Gibbs posterior of $\theta$ concentrates toward $\theta_{0}$ and enjoys asymptotic normality under mild conditions (Miller, 2021).
@@
The second methodological gap relates to the modeling front: continuity of the Gibbs posterior distribution often yields a mismatch to constraint conditions on $\hat{\theta}$ except on a set of measure zero. To illustrate, consider the Bayesian lasso (Park and Casella, 2008), which can be viewed as the Gibbs posterior under the lasso loss. Though this promotes a sparse estimate $\hat{\theta}$, under its posterior distribution $\theta$ is non-sparse almost everywhere. A similar problem arises in a Gibbs posterior approach to support vector machines. The maximum-margin hyperplane has zero posterior measure, which may partially explain why studies from this view have focused on point estimation (Polson and Scott, 2011), and motivates our approach in seeking a more natural quantification of the associated uncertainty. Beyond this incongruence between $\hat{\theta}$ and the samples from $\Pi(\theta\mid y)$, invariance to changes in $g(\theta,y)$ presents another consideration. To obtain an estimate $\hat{\theta}$ residing in a constrained or low dimensional space, it is common practice in optimization to employ an alternative $\tilde{g}(\theta,y)$ that has superior computational properties. For example, $\tilde{g}$ can be convex, unconstrained, or non-combinatorial, under the condition that ${\arg\min}_{\theta}\tilde{g}(\theta,y)={\arg\min}_{\theta}g(\theta,y)$—that is, the two distinct loss functions touch at the minima. This invariance at the optimum is routinely exploited in methods such as convex relaxation, variable splitting, proximal methods, and majorization-minimization (Polson et al., 2015; Zheng and Aravkin, 2020; Landeros et al., 2023). However, the Gibbs posterior does not enjoy such an invariance, as the distribution $\Pi(\theta\mid y)$ changes whenever $g$ changes.
@@
These issues lead us to a marked departure from existing approaches. Rather than treating $\theta$ as a high-dimensional random variable, we model $\theta=(z,\lambda)$ with only $\lambda$ as a parameter with a corresponding prior distribution. The argument $z$ is instead treated as a latent variable that is deterministic conditional on $y$ and $\lambda$, though importantly it remains a stochastic quantity when conditioned on $y$ alone. As we will demonstrate in the article, this effectively reduces the dimension of $\theta$ to near that of $\lambda$, simultaneously addressing both issues surveyed above. Specifying $z$ as the solution of an optimization subproblem allows us to retain transparent constraint conditions such as low rank, low cardinality, or combinatorial constraints.
It is natural to ask whether such an approach is consistent with Bayesian methodology in the sense that there exists a valid generative model corresponding to a likelihood that depends only on $\lambda$. This article answers this question affirmatively. We begin with a set of profile likelihoods that partially maximize a joint likelihood $L(y;z,\lambda)$ over $z$, showing that each corresponds to another common likelihood where the data are modeled dependently. We then establish the theoretical result that under mild conditions, the $\sqrt{n}$-adjusted posterior distribution of the parameter under our framework converges asymptotically to the same normal limit as canonical posteriors marginalized over non-deterministic latent variables.
@@
This contribution is closely related to prior work by Polson and Scott (2016), which discovers a hierarchical duality: the scale mixture of univariate exponential or location-scale mixture of normal is proportional to another (potentially intractable) density maximized over a univariate latent variable. This perspective inspires efficient new algorithms for producing point estimates. Despite some similarities in the univariate setting, our method applies generally to multivariate problems and to settings where the latent variables may exhibit dependence. In other related work, Lee et al. (2005) interpret the profile likelihood as resulting from an empirical prior. A key difference is that our proposed framework $L(y;z,\lambda)$ can lead to a fully Bayesian method, where the latent variable $z$ characterizes the latent dependency among the data. The source code can be found on https://github.com/Zeng-Cheng/bridged_posterior_code_for_paper.
@@ 2.1 Augmented Likelihood with Conditional Optimization
To provide background, we first review the canonical likelihood involving latent variables, taking the form
@@
in which we refer to $\lambda\in\mathbb{R}^{d}$ as parameter, and $z\in\mathbb{R}^{p}$ as latent variable. Here $\Pi_{\mathcal{L}}$ denotes the marginal latent variable distribution for $z$. Since $z$ could be associated with a continuous, discrete, or degenerate distribution, we use the integration with respect to a probability measure notation $\int f(z)\mu(\textup{d}z)$, in which $z$ with distribution $z\sim\mu$ is the one that we integrate over. The joint distribution $L(y,z;\lambda)$ is also known as an “augmented likelihood” (Tanner and Wong, 1987; Van Dyk and Meng, 2001). Examples abound in statistics: for instance, augmented likelihoods are used in characterizing dependence among discrete $y$ via a correlated normal latent variable $z$ (Wolfinger, 1993; Rue et al., 2009), or model-based clustering on grouping data $y$ via a latent discrete label $z$ (Blei et al., 2003; Fraley and Raftery, 2002).
We now consider a special case when given $y$ and $\lambda$:
@@
For simplicity of exposition, from here we focus on the case where $z$ is the unique minimizer. This encompasses a large class of models and is satisfied whenever $g(\zeta,y;\lambda)$ is strictly convex in $\zeta$ for every $(y,\lambda)$. Though $z$ is conditionally deterministic form, note that when we do not condition on $y$, $z$ remains randomly distributed under $\Pi_{\mathcal{L}}(z;\lambda)$. This suggests a generative view according to (1): we have
$z\sim\Pi_{\mathcal{L}}(z;\lambda);\qquad$ $y\mid z,\lambda\sim L(y\in\mathcal{Y}_{\lambda,z}\mid z,\lambda),\quad\text{ where }\mathcal{Y}_{\lambda,z}=\big{\{}y:{\min}_{\zeta}g(\zeta,y;\lambda)=g(z,y;\lambda)\big{\}}.$
That is, $y$ is generated under the constraint given by $z$.
Remark 1.
To clarify, the latent $z$ may have varying dimension $p$ and $\Pi_{\mathcal{L}}$ according to the sample size $n$.
For concreteness, we present two illustrative examples based on the profile likelihood. Profile likelihoods have a frequentist origin, motivated by the convenience of testing or constructing confidence intervals for a parameter of interest $\lambda$, in the presence of other parameters $\zeta$, often called “nuisance parameters”. There is a long-standing debate on whether using a profile likelihood leads to a coherent Bayes’ procedure (Lee et al., 2005; Cheng and Kosorok, 2009; Evans, 2016; Maclaren, 2018). Using the above, we can now view the profile likelihood as a special case of (2), taking $g(\zeta,y;\lambda)=-\log L(y,\zeta;\lambda)$.
@@
Consider linear regression with $y\in\mathbb{R}^{n}$, $X\in\mathbb{R}^{n\times d}$, $\lambda\in\mathbb{R}^{d}$, $z>0$, and some constant $v>0$:
@@
The first line has the same form as a likelihood with normal errors, with the variance $z$ regularized by an Inverse-Gamma($v/2$, $v/2$). Now, instead of marginalizing out $z$, we may maximize $\log L(y,\zeta;\lambda)$ over $\zeta$ to obtain the closed-form $z=(v+\|y-X\lambda\|^{2})/(v+n+2)$. Therefore, we have
$\Pi_{\mathcal{L}}(z;\lambda)\propto z^{-(n+v+2)/2};\quad$ $L(y\mid z,\lambda)\propto\exp\bigg{(}-\frac{\|y-X\lambda\|^{2}}{2z}\bigg{)}\mathbb{1}\biggl{\{}\sum_{i=1}^{n}\bigl{(}y_{i}-x_{i}^{\top}\lambda\bigr{)}^{2}=(v+n+2)z-v\biggr{\}}.$
In particular, the indicator above imposes a quadratic equality constraint on $y$. Upon substituting an expression in $y$ for $z$, we obtain a marginal density
@@
$L(y,z;\lambda)\propto\exp\biggl{(}-\frac{\|y-Cz\|^{2}}{2\sigma^{2}}\biggr{)}\exp\biggl{(}-\frac{1}{2}z^{\top}G^{-1}z\biggr{)};$
$z={\arg\min}_{\zeta}\{-\log L(y,\zeta;\lambda)\}.$
The first line has the same form as a likelihood with $z$ regularized by multivariate normal distribution $\text{N}(0,G)$. Here, minimization yields $z=(C^{\top}C/\sigma^{2}+G^{-1})^{-1}C^{\top}y/\sigma^{2}=\{G-GC^{\top}(I\sigma^{2}+CGC^{\top})^{-1}CG\}C^{\top}y/\sigma^{2}$. Therefore,
$\Pi_{\mathcal{L}}(z;\lambda)\propto\exp\left[-\frac{1}{2}z^{\top}\{GC^{\top}(I\sigma^{2}+CGC^{\top})^{-1}CG\}^{-1}z\right];$
@@
In this example, the indicator puts an affine equality constraint on $y$. On the other hand, the marginal of $y$ is
@@
which corresponds to a multivariate normal $\text{N}\bigl{(}0,I\sigma^{2}+CGC^{\top}\bigr{)}$.
@@
The profile likelihood-based models are an important sub-class that we will primarily focus on. On the other hand, in general, the loss function $g$ does not have to be the negative likelihood, and $z$ does not have to be available in closed form. We can still specify the joint likelihood, by including an optimization problem in the equality constraint (2).

### v2

# Bridged Posterior: Optimization, Profile Likelihood and a New Approach to Generalized Bayes

@@ Abstract
Optimization is widely used in statistics, and often efficiently delivers point estimates on useful spaces involving structural constraints or combinatorial structure. To quantify uncertainty, Gibbs posterior exponentiates the negative loss function to form a posterior density. Nevertheless, Gibbs posteriors are supported in high-dimensional spaces, and do not inherit the computational efficiency or constraint formulations from optimization. In this article, we explore a new generalized Bayes approach, viewing the likelihood as a function of data, parameters, and latent variables conditionally determined by an optimization sub-problem. Marginally, the latent variable given the data remains stochastic, and is characterized by its posterior distribution. This framework, coined bridged posterior, conforms to the Bayesian paradigm. Besides providing a novel generative model, we obtain a positively surprising theoretical finding that under mild conditions, the $\sqrt{n}$-adjusted posterior distribution of the parameters under our model converges to the same normal distribution as that of the canonical integrated posterior. Therefore, our result formally dispels a long-held belief that partial optimization of latent variables may lead to underestimation of parameter uncertainty. We demonstrate the practical advantages of our approach under several settings, including maximum-margin classification, latent normal models, and harmonization of multiple networks.
KEY WORDS: Constrained conditional probability; Data augmentation; Duality; Latent variable model; Parametric and semi-parametric Bernstein–von Mises.
@@ 1 Introduction
The generalized Bayes approach is becoming increasingly popular due to its potential advantages in model simplicity and robustness. A generalized posterior can be specified based on partial information from the data, or via a loss function that characterizes an inferential summary of the data. This appeals when the likelihood is inaccessible or intractable; there is a well-established literature on partial information settings including methods based on composite likelihood (Lindsay, 1988; Varin et al., 2011), partial likelihood (Sinha et al., 2003; Dunson and Taylor, 2005), pairwise likelihood (Jensen and Künsch, 1994), and others. Recently, there has been a burgeoning interest in loss-based Bayesian models, including works involving classification loss (Polson and Scott, 2011) or distance-based losses (Duan and Dunson, 2021; Rigon et al., 2023; Natarajan et al., 2024). Loss-based generalized Bayes models typically use a probability distribution called the Gibbs posterior (Jiang and Tanner, 2008), taking the form: $\Pi(\theta\mid y)\propto\exp\{-g(\theta,y)\},$ where $g(\theta,y)$ is some loss function with $y$ the data and $\theta$ the parameter.
@@
There is a vast generalized Bayes literature using the Gibbs posterior for explicit model weighting, with $g$ chosen according to utility functions such as predictive accuracy (Lavine et al., 2021; Tallman and West, 2024), scoring rule (Gneiting and Raftery, 2007; Dawid and Musio, 2015), fairness metrics (Chakraborty et al., 2024) or summary statistics-based divergence (Frazier and Drovandi, 2021; Frazier et al., 2023). Such an approach also lends itself to modular descriptions of data (Jacob et al., 2017), and can guard against model misspecification (Nott et al., 2023). With connections to these methods, our focus is on the case when one wants to adopt a loss $g$ from the optimization literature for statistical modeling, while needing to quantify uncertainty beyond point estimates.
@@
The point estimate $\hat{\theta}=\arg\min_{\theta}g(\theta,y)$ can often be efficiently computed using an iterative optimization algorithm, even under a wide range of constraints. For example, convex clustering and its variants (Tan and Witten, 2015; Chi and Lange, 2015; Chakraborty and Xu, 2023) use $g(\theta,y)=(1/2)\sum_{i=1}^{n}\|y_{i}-\theta_{i}\|^{2}_{2}+\lambda\sum_{(i,j):i<j}\|\theta_{i}-\theta_{j}\|_{2}$ for data $y_{i}\in\mathbb{R}^{p}$, location parameter $\theta_{i}\in\mathbb{R}^{p}$, and tuning constant $\lambda>0$. This can be understood as a relaxation of hierarchical clustering; in place of a combinatorial constraint, the penalty term encourages most of the $L_{2}$-norms $\|\hat{\theta}_{i}-\hat{\theta}_{j}\|_{2}$ to be zero, promoting cluster structure via a small number of unique $\hat{\theta}_{i}$ at the solution. The estimate $\hat{\theta}$ can be obtained using convex, continuous optimization. A popular combinatorial alternative makes use of the $k$-means loss (MacQueen, 1967) toward clustering, $g\{(c_{1},\ldots,c_{n}),y\}=\sum_{k=1}^{K}\sum_{(i<j):c_{i}=c_{j}=k}\|y_{i}-y_{j}\|^{2}_{2}/n_{k}$, with $c_{i}\in\{1,\ldots,K\}$ where the discrete cluster assignment label $c_{i}=k$ if $\theta_{k}$ is the nearest centroid to $y_{i}$, $n_{k}=\sum_{i}1(c_{i}=k)$, and $\hat{\theta}_{i}=\sum_{i:c_{i}=k}y_{i}/n_{k}$. Here too, iterative algorithms can improve performance and avoid local minima using continuous optimization techniques (Xu and Lange, 2019). Recently, Rigon et al. (2023) form a Gibbs posterior using this loss toward quantifying the uncertainty of $c_{i}$, which shows robustness to the distributional asymmetry. Several recent works import ideas from optimization to account for constraints within a Bayesian framework (Duan et al., 2020; Presman and Xu, 2023; Zhou et al., 2024).
@@
Under the exponential negative transformation, the Gibbs posterior distribution concentrates near the posterior mode. This induces variability around the point estimate and, in turn, enables uncertainty quantification. How to interpret this uncertainty is not immediately obvious, and one may question the authenticity of inferential procedures such as hypotheses tests or intervals based on such a posterior which may not derive from a generative likelihood model. There are several works that provide justification in the large $n$ regime. First, Gibbs posteriors admit a coherent update scheme for $\theta$ toward minimizing the expected loss $\int g(\theta,y)\mathcal{F}(\textup{d}y;\theta)$, where $\mathcal{F}$ denotes the true data generating distribution (Bissiri et al., 2016). Second, if the Gibbs posterior density is proportional to a composite likelihood, such as the conditional density under some insufficient statistic (Lewis et al., 2021) derived from a full likelihood $F(y;\theta_{0})$, then the Gibbs posterior of $\theta$ concentrates toward $\theta_{0}$ and enjoys asymptotic normality under mild conditions (Miller, 2021).
@@
The second methodological gap relates to the modeling front: continuity of the Gibbs posterior distribution often yields a mismatch to constraint conditions on $\hat{\theta}$ except on a set of measure zero. To illustrate, consider the Bayesian lasso (Park and Casella, 2008), which can be viewed as the Gibbs posterior using the lasso loss. Though this promotes a sparse estimate $\hat{\theta}$, under its posterior distribution $\theta$ is non-sparse almost everywhere. A similar problem arises in a Gibbs posterior approach to support vector machines. The maximum-margin hyperplane has zero posterior measure, which may partially explain why studies from this view have focused on point estimation (Polson and Scott, 2011), and motivates our approach in seeking a more natural quantification of the associated uncertainty. Beyond this incongruence between $\hat{\theta}$ and the samples from $\Pi(\theta\mid y)$, invariance to changes in $g(\theta,y)$ presents another consideration. To obtain an estimate $\hat{\theta}$ residing in a constrained or low-dimensional space, it is common practice in optimization to employ an alternative $\tilde{g}(\theta,y)$ that has superior computational properties. For example, $\tilde{g}$ can be convex, unconstrained, or non-combinatorial, under the condition that ${\arg\min}_{\theta}\tilde{g}(\theta,y)={\arg\min}_{\theta}g(\theta,y)$—that is, the two distinct loss functions touch at the minima. This invariance at the optimum is routinely exploited in methods such as convex relaxation, variable splitting, proximal methods, and majorization-minimization (Polson et al., 2015; Zheng and Aravkin, 2020; Landeros et al., 2023). However, the Gibbs posterior does not enjoy such an invariance, as the distribution $\Pi(\theta\mid y)$ changes whenever $g$ changes.
@@
These issues lead us to take a marked departure from existing approaches. Rather than treating $\theta$ as a high-dimensional random variable, we model $\theta=(z,\lambda)$ with only $\lambda$ as a parameter with a corresponding prior distribution. The argument $z$ is instead treated as a latent variable that is deterministic conditional on $y$ and $\lambda$, though importantly it remains a stochastic quantity when conditioned on $y$ alone. As we will demonstrate in the article, this effectively reduces the dimension of $\theta$ to nearly that of $\lambda$, simultaneously addressing both issues surveyed above. Specifying $z$ as the solution of an optimization subproblem allows us to retain transparent constraint conditions such as low rank, low cardinality, or combinatorial constraints.
@@
It is natural to ask whether such an approach is consistent with Bayesian methodology, that there exists a valid generative model corresponding to a likelihood that depends only on $\lambda$. This article answers this question affirmatively. We begin with a set of profile likelihoods that partially maximize a joint likelihood $L(y;z,\lambda)$ over $z$, showing that each corresponds to another common likelihood where the data are modeled dependently. We then establish the theoretical result that under mild conditions, the $\sqrt{n}$-adjusted posterior distribution of the parameter under our framework converges asymptotically to the same normal limit as canonical posteriors marginalized over non-deterministic latent variables. This contribution is closely related to prior work by Polson and Scott (2016), which discovers a hierarchical duality: the scale mixture of univariate exponential or location-scale mixture of normal is proportional to another (potentially intractable) density maximized over a univariate latent variable. This perspective inspires efficient new algorithms for producing point estimates. Despite some similarities in the univariate setting, our method applies generally to multivariate problems and to settings where the latent variables may exhibit dependence. In other related work, Lee et al. (2005) interpret the profile likelihood as resulting from an empirical prior. A key difference is that our proposed framework $L(y;z,\lambda)$ can lead to a fully Bayesian method, where the latent variable $z$ characterizes the latent dependency among the data. The source code can be found on https://github.com/Zeng-Cheng/bridged_posterior_code_for_paper.
@@ 2.1 Augmented Likelihood with Conditional Optimization
To provide background, we first review the canonical likelihood involving latent variables, which takes the form
@@
in which we refer to $\lambda\in\mathbb{R}^{d}$ as the parameter, and $z\in\mathbb{R}^{p}$ as the latent variable. Here $\Pi_{\mathcal{L}}$ denotes the marginal latent variable distribution for $z$. Since $z$ could be associated with a continuous, discrete, or degenerate distribution, we use the integration with respect to a probability measure notation $\int f(z)\mu(\textup{d}z)$, in which $z$ with distribution $z\sim\mu$ is the one that we integrate over. The joint distribution $L(y,z;\lambda)$ is also known as an augmented likelihood (Tanner and Wong, 1987; Van Dyk and Meng, 2001). Examples abound in statistics: for instance, augmented likelihoods are used in characterizing dependence among discrete $y$ via a correlated normal latent variable $z$ (Wolfinger, 1993; Rue et al., 2009), or model-based clustering on grouping data $y$ via a latent discrete label $z$ (Blei et al., 2003; Fraley and Raftery, 2002). We now consider a special case when given $y$ and $\lambda$:
@@
For simplicity of exposition, from here we focus on the case where $z$ is the unique minimizer. This encompasses a large class of models and is satisfied whenever $g(\zeta,y;\lambda)$ is strictly convex in $\zeta$ for every $(y,\lambda)$. Though $z$ is conditionally deterministic, it is important to note that when we do not condition on $y$, $z$ remains randomly distributed under $\Pi_{\mathcal{L}}(z;\lambda)$. This suggests a generative view according to (1): we have
$z\sim\Pi_{\mathcal{L}}(z;\lambda);\qquad y\mid z,\lambda\sim L(y\in\mathcal{Y}_{\lambda,z}\mid z,\lambda),$ (3)
[…]
That is, $y$ is generated under the constraint given by $z$. This formulation allows the latent $z$ to have varying dimension $p$ and $\Pi_{\mathcal{L}}$ according to the sample size $n$.
For concreteness, we present two illustrative examples based on the profile likelihood. Profile likelihoods have frequentist origins, motivated by the convenience of testing or constructing confidence intervals for a parameter of interest $\lambda$ in the presence of other nuisance parameters $\zeta$. There is a long-standing debate on whether using a profile likelihood leads to a coherent Bayes’ procedure (Lee et al., 2005; Cheng and Kosorok, 2009; Evans, 2016; Maclaren, 2018). Using the above, we can now view the profile likelihood as a special case of (2), taking $g(\zeta,y;\lambda)=-\log L(y,\zeta;\lambda)$.
@@
Consider linear regression with $y\in\mathbb{R}^{n}$, $X\in\mathbb{R}^{n\times d}$, $\lambda\in\mathbb{R}^{d}$, $z>0$, $v>0$:
@@
The first line has the same form as a likelihood with normal errors, with the variance $z$ regularized by an Inverse-Gamma($v/2$, $v/2$). Instead of marginalizing out $z$, we maximize $\log L(y,\zeta;\lambda)$ over $\zeta$ to obtain $z=(v+\|y-X\lambda\|^{2})/(v+n+2)$. Therefore, we have
$\Pi_{\mathcal{L}}(z;\lambda)\propto z^{-(n+v+2)/2};$
$L(y\mid z,\lambda)\propto\exp\bigg{(}-\frac{\|y-X\lambda\|^{2}}{2z}\bigg{)}\mathbb{1}\biggl{\{}\sum_{i=1}^{n}\bigl{(}y_{i}-x_{i}^{\top}\lambda\bigr{)}^{2}=(v+n+2)z-v\biggr{\}}.$
In particular, the indicator above imposes a conditional constraint $\mathcal{Y}_{\lambda,z}$ on $y$ given $z$, which corresponds to a ball centered at $X\lambda$ with radius $\sqrt{(v+n+2)z-v}$. Upon substituting an expression in $y$ for $z$, we obtain a marginal density
@@
$L(y,z;\lambda)\propto\exp\biggl{(}-\frac{\|y-Cz\|^{2}}{2\sigma^{2}}\biggr{)}\exp\biggl{(}-\frac{1}{2}z^{\top}G^{-1}z\biggr{)};$ $z={\arg\min}_{\zeta}\{-\log L(y,\zeta;\lambda)\}.$
The first part has the same form as a likelihood with $z$ regularized by a multivariate normal distribution $\text{N}(0,G)$. Here, minimization yields $z=(C^{\top}C/\sigma^{2}+G^{-1})^{-1}C^{\top}y/\sigma^{2}=\{G-GC^{\top}(I\sigma^{2}+CGC^{\top})^{-1}CG\}C^{\top}y/\sigma^{2}$. Therefore,
$\Pi_{\mathcal{L}}(z;\lambda)\propto\exp\biggl{[}-\frac{1}{2}z^{\top}\{GC^{\top}(I\sigma^{2}+CGC^{\top})^{-1}CG\}^{-1}z\biggr{]};$
@@
The indicator again imposes a conditional constraint $\mathcal{Y}_{\lambda,z}$ on $y$ given $z$, in this case corresponding to an affine subspace of $\mathbb{R}^{\tilde{p}}$ with dimension $\tilde{p}-p$. The marginal of $y$ is
@@
corresponding to a multivariate normal $\text{N}\bigl{(}0,I\sigma^{2}+CGC^{\top}\bigr{)}$.
@@
The profile likelihood-based models are an important sub-class that we will primarily focus on. Nevertheless, in general, the loss function $g$ does not have to be the negative likelihood, and $z$ does not have to be available in closed form. We can still specify the joint likelihood, by including an optimization problem in the equality constraint (2).
@@
Despite the connection with the canonical full likelihood, in which $\lambda$ would be marginalized, the bridged posterior should be interpreted as a distinct generative model. Specifically, based on (LABEL:eq:constrained_view) under the bridged posterior, $y$ has a distribution supported on a constrained space $\mathcal{Y}_{\lambda,z}$ given $z$. In contrast, the canonical full likelihood typically does not feature such a conditional constraint.